OpenStudy (samigupta8):

a monkey while trying to reach the top of a pole of height 12 metres takes every time a jump of 2 metres but slips 1 metre while holding the pole.the number of jumps required to reach the top of the ople is

OpenStudy (anonymous):

The number of jumps required to reach the top of the pole is what?

OpenStudy (samigupta8):

exactly that's my question

OpenStudy (swissgirl):

Well so basically in other terms we can say that the monkey only jumps 1 meter every jump

OpenStudy (swissgirl):

Since every jump he moves up 2 meters but slips 1 meter which is 2-1=1

OpenStudy (anonymous):

Oh haha I read that wrong!..Im sorry.. but swiss girl is right.

OpenStudy (swissgirl):

So we can model our equation as follow 2x-x=12 Or (2-1)x=12 and then solve for x

OpenStudy (swissgirl):

or x=12

OpenStudy (anonymous):

It takes 11 jumps. He won't slip from 12 to 11 on the last jump because he's already at the top. You are basically counting the elements in the following: 2, 3, 4, . . . , 12 which is 11 each element is the position before the slip but after the jump.

OpenStudy (jdoe0001):

well, why would a monkey be jumping on a bare pole anyway? is not like it has figs or bananas, so really on every jump the monkey is wasting 2 meters plus momentum of eneergy :P

OpenStudy (swissgirl):

Hmmm I think tcarroll is correct

OpenStudy (samigupta8):

u r right tcarrol.....

OpenStudy (anonymous):

thanks much. That was a fun problem! That monkey better get a better grip!